### v2

@@ 5.3 Expanding Our Cluster Analysis to Task Embeddings
Until now, our analysis has primarily focused on the embeddings of models. However, during PetriNet2Vec training, we also learn task embeddings, meaning that each task depicted in Figure 4a is also mapped to an 8-dimensional embedding vector. Employing HDBSCAN on these task embeddings revealed the emergence of five clusters. Figure 8a displays the Silhouette plot, while Figure 8b illustrates the 2-dimensional projection of these task vectors. Unlike the clusters of models, the clusters of tasks exhibit lower quality, and some tasks were not associated with any cluster, indicated as white dots (or cluster noises). This occurred because the cosine similarity with their closest neighbor was considered too large by the HDBSCAN algorithm. Nonetheless, they still exhibit some degree of cohesion, and most importantly, they play a significant role during the training of model embeddings, granting PetriNet2Vec greater flexibility in accommodating those model embeddings.
Upon examining the scatter plot in Figure 8a, we observed a tendency for tasks sharing OR-splits, such as $t_{2}$, $t_{39}$, and $t_{40}$, to be grouped together, likely owing to their resemblance to text synonyms in doc2vec. By analogy, tasks preceding an OR-split create alternate paths within the models, enabling the process to achieve the same subsequent state through diverse routes. Intriguingly, some white tasks are frequently subject to substitution by black boxes when the $D=1$ rule is applied. As these tasks are replaced, their occurrence diminishes in PNML files, consequently reducing their sampling during training. Moreover, infrequent tasks exist in models that are not necessarily substituted by black boxes. Analogously, these tasks mimic the behavior of less frequent words in doc2vec. This becomes apparent when contrasting the histogram in Figure 4a with Figure 8b, where less frequent tasks are predominantly denoted as white dots. Nonetheless, while other analogies may be drawn, it is crucial to acknowledge that certain cluster artifacts may arise due to the inherent limitations of HDBSCAN.
@@
(a) Silhouette from task embeddings
(b) UMAP from task embeddings
@@ 5.4 Examples of Downstream Tasks
Addressing the complexity and significance of querying business process models provides invaluable insights for large enterprises managing numerous processes. It enables them to pinpoint structurally similar processes, facilitating informed decisions such as merging or replacing process models. Moreover, this methodology can be extended to identifying similar tasks within processes. Figure 11 visually demonstrates an investigation of this nature, illustrating the similarity between each task embedding and each model embedding. In this figure, the models are ordered according to the clusters they formed during the application of HDBSCAN. Notably, certain task embeddings closely align with all model embeddings within the same cluster, unveiling compelling patterns. Nevertheless, interpreting task-model embeddings demands caution. Although they share the same embedding dimensions, they are different vector spaces. Consequently, the resulting score would not directly translate to a strict interpretation like model-to-model similarity. Nevertheless, it still offers valuable insights into the task’s relevance to the model’s knowledge.
@@
Our second example of a downstream task for model embeddings is classifying processes based on formation rules. In this case, we train a classifier using the embedding vectors as features and the rules described in Section 3.3 as labels. Six k-NN classifiers were trained using 5-fold cross-validation and cosine distance. The results can be found in the confusion matrices shown in Figure 12. In this figure, we can observe that some properties of the models, mainly those that achieved higher accuracy, were effectively incorporated by our methodology when creating the embeddings, such as loops, OR-constructs, invisible tasks, and duplicated tasks. However, bypass connections (representing long-term tasks) and optional tasks were more challenging to recognize from the learned embedding, although the results can still be considered satisfactory. This limitation probably arises from the fact that our methodology only considers pairs of consecutive tasks, neglecting, for example, long-term dependencies.

### v3

@@ 5.3 Expanding Our Cluster Analysis to Task Embeddings
Until now, our analysis has primarily focused on the embeddings of models. However, during PetriNet2Vec training, we also learn task embeddings, meaning that each task depicted in Figure 4a is also mapped to an 8-dimensional embedding vector. Employing HDBSCAN on these task embeddings revealed the emergence of five clusters. Figure 8a displays the Silhouette plot, while Figure 8b illustrates the 2-dimensional projection of these task vectors.
Unlike the clusters of models, the clusters of tasks exhibit lower quality, and some tasks were not associated with any cluster, indicated as white dots (or cluster noises). This occurred because the cosine similarity with their closest neighbor was considered too large by the HDBSCAN algorithm. Nonetheless, they still exhibit some degree of cohesion, and most importantly, they play a significant role during the training of model embeddings, granting PetriNet2Vec greater flexibility in accommodating those model embeddings.
Upon examining the scatter plot in Figure 8a, we observed a tendency for tasks sharing OR-splits, such as $t_{2}$, $t_{39}$, and $t_{40}$, to be grouped together, likely owing to their resemblance to text synonyms in doc2vec. By analogy, tasks preceding an OR-split create alternate paths within the models, enabling the process to achieve the same subsequent state through diverse routes.
Intriguingly, some white tasks are frequently subject to substitution by black boxes when the $D=1$ rule is applied. As these tasks are replaced, their occurrence diminishes in PNML files, consequently reducing their sampling during training. Moreover, infrequent tasks exist in models that are not necessarily substituted by black boxes. Analogously, these tasks mimic the behavior of less frequent words in doc2vec. This becomes apparent when contrasting the histogram in Figure 4a with Figure 8b, where less frequent tasks are predominantly denoted as white dots. Nonetheless, while other analogies may be drawn, it is crucial to acknowledge that certain cluster artifacts may arise due to the inherent limitations of HDBSCAN.
@@
(a) Silhouette from task embeddings.
(b) UMAP from task embeddings.
@@ 5.4 Examples of Downstream Tasks
Addressing the complexity and significance of querying business process models provides invaluable insights for large enterprises managing numerous processes. It enables them to pinpoint structurally similar processes, facilitating informed decisions such as merging or replacing process models. Moreover, this methodology can be extended to identifying similar tasks within processes. Figure 11 visually demonstrates an investigation of this nature, illustrating the similarity between each task embedding and each model embedding. In this figure, the models are ordered according to the clusters they formed during the application of HDBSCAN. Notably, certain task embeddings closely align with all model embeddings within the same cluster, unveiling compelling patterns.
Nevertheless, interpreting task-model embeddings demands caution. Although they share the same embedding dimensions, they are different vector spaces. Consequently, the resulting score would not directly translate to a strict interpretation like model-to-model similarity. Nevertheless, it still offers valuable insights into the task’s relevance to the model’s knowledge.
@@
Our second example of a downstream task for model embeddings is classifying processes based on formation rules. In this case, we train a classifier using the embedding vectors as features and the rules described in Section 3.3 as labels. Six k-NN classifiers were trained using 5-fold cross-validation and cosine distance.
The results can be found in the confusion matrices shown in Figure 12. In this figure, we can observe that some properties of the models, mainly those that achieved higher accuracy, were effectively incorporated by our methodology when creating the embeddings, such as loops, OR-constructs, invisible tasks, and duplicated tasks. However, bypass connections (representing long-term tasks) and optional tasks were more challenging to recognize from the learned embedding, although the results can still be considered satisfactory. This limitation probably arises from the fact that our methodology only considers pairs of consecutive tasks, neglecting, for example, long-term dependencies.
@@ 6 Conclusions
One limitation we observed during the clustering experiments was that the length of the context window negatively impacts clustering performance. Addressing this issue could enable the characterization of long-range paths within the Petri nets. Furthermore, the PetriNet2Vec model needs to be retrained when applied to new process spaces that utilize a different set of tasks. This retraining requirement poses a challenge in adapting the model to novel process domains without additional training data and computational resources.